OpenStudy (anonymous):

Explain how to convert f(x) into the general, vertex form of the equation. Use complete sentences. x2+6x+1

hero (hero):

Find the vertex of x^2 + 6x + 1. Then insert a = 1 and vertex (h,k) into the equation y = a(x - h)^2 + k

OpenStudy (anonymous):

I don't understand may you explain

hero (hero):

That's the easy way to do it, but since you probably have to complete the square, you should do it that way instead.

OpenStudy (cwrw238):

to complete the square take half the coefficient of x ( half of 6 = 3) and write (x + 3)^2 to make this equal to x^2 + 6x we need to subtract 3^2 :- x^2 + 6x = (x + 3)2 - 9 now we add the 1 x^2 + 6x + 1 = (x+3)^2 - 9 + 1 = (x +3)^2 - 8 which is the vertex form the vertex is at the point (-3,-8) (compare with Hero's formula)

OpenStudy (anonymous):

and how do you find the solutions

OpenStudy (cwrw238):

you have it there - you wanted the vertex form

OpenStudy (cwrw238):

do you want the zeroes of the function as well?

OpenStudy (anonymous):

yes

OpenStudy (cwrw238):

(x +3)^2 - 8 = 0 (x +3)^2 = 8 now take the square root of each side of the equation (remember the positive and negative square root of 8)

OpenStudy (cwrw238):

the rest is easy